OpenStudy (ashsmashum1290):

A circle has a radius of 3 cm. What happens to the area if the radius is doubled to 6 cm. A.It is doubled B.It is tripled C.It is quadrupled D.It stays the same

OpenStudy (ashsmashum1290):

@Hero please help

OpenStudy (texaschic101):

lets find out... area of a circle = (pi) * r^2 and if the radius is 3.....then area is : 3.14(3^2) = 3.14 * 9 = 28.26 and if the radius is 6 area = 3.14(6^2) = 3.14 * 36 = 113.04 28.26 * 4 = 113.04.....the area is quadrupled

OpenStudy (texaschic101):

quadrupled means multiply by 4

OpenStudy (ashsmashum1290):

so its C?

OpenStudy (ashsmashum1290):

@texaschic101

OpenStudy (anonymous):

C because the are with the radius of 3 is 9pi, while with the radius of 6 the area is 36pi

OpenStudy (anonymous):

36/9=4 quadrupled

OpenStudy (ashsmashum1290):

okay thanks

OpenStudy (anonymous):

No medal?